LUCIFER Technical Observing, prep for Partner Science: UT 2010 January 03

Observer: DThompson, MPedani (LBTO)
Telescope Operator: WWack
Instrument Support: DOfficer, WSeifert (on call)
Telescope Support: JUrban
SW support: TLeibold (on call)

Summary

  • Twilight flats in Brg, H2, J, H, Ks
  • J data on Garnavich SDT project (guiding tests)
  • ptcorr script tested - pointing correction now takes only seconds
  • open-loop tracking tests
  • made and implemented new collimation model
  • Took some field aberration measurements
  • Observed 52 pointing stars
    • Model has 3.7 arcsec rms, implemented as default on 20100104
    • LUCIFERSX.20100103.ptmod
  • tested updatePointingReference command
  • Took Martini's JHKs imaging data
  • Only significant problem was an OSS hang - 45 minutes lost

Preparations

- Fix M1 M2 Y global offset discrepancies
  • JMH Fixed the global offset problems on 27-DEC-2009.
  • JMH adjusted LUCIFERSX.20091227.ptmod to have the new IA value of -260. Note that this change essentially puts us back to the 24-Sep-2009 version of the LUCIFER pointing model.

Details

~22:45 LUCIFER UI started. No issues. Telescope is in the process of configuration change and rebalance for LUCIFER operations.

Files: luci.20100103.NNNN.fits

23:50 Telescope ready for twilight flats at Mirror is +2C relative to outside temp.

0:00 Start twilight flats:
Br_gam 2 sec 0004-0013 flux~4700 counts
Br_gam 5 sec 0014-0023 flux~11000 counts
H2 2 sec 0024-0033 flux~6000 counts

00:35 Br_gam 2 sec 0034-0043 flux~2000 counts
00:39 Br_gam 2 sec 0044-0053 flux~1100 counts
Br_gam 5 sec 0054-0063 flux~2500 counts
H2 2 sec 0064-0073 flux~800 counts

00:45 Ks 5 sec 0074-0083
H 2 sec 0084-0093
J 2 sec 0094-0103
Ks 5 sec 0104-0113 flux~1700 counts
H 2 sec 0114-0123 flux~1200 counts
J 2 sec 0124-0133 flux~1100 counts

01:03 Correct initial collimation/pointing with IE/CA, test ptcorr script
Pointing correction. The ptcorr script worked (now ptmodifies take ~5 seconds). But this is where OSS hung. Initial collimation has obvious coma and considerable defocus.

01:12 LUCIFER Telescope Control got stuck in mid-commit. Initialize did not recover. Changing to wait for collimation off did not recover. TCS appears to be stuck somewhere between wait for elevation complete and wait for telescope started. Called Torsten at 01:18. He advised to call Chris. 01:19 called Chris. He points out that the IIF GUI looks stuck at "Collimating M2 (PSF)". It looks like OSS was stuck trying to collimate M2. Stop OSS did not work, kill OSS took a while to do so. LUCIFER UI cam back when the preset failed. OSS was the problem. Lost 45 minutes.

01:45
Run short dither dice5 pattern(s) to check that things still work as in May/June
Check guide star FWHM/rms vs LUCIFER image quality (closed loop DIQ)
Both tests can be done on an SDT target: execute garnavich_osurc.scriptJ. Corrected the END_OSERVING_SETUP and script tested OK. Wait for collimation was still off, but the initial guide images look good and are ~0.6 arcsec FWHM (initial collimation cycle below 600nm). Files 0139-0161. Seeing ~5.5 pixels in J (0.66 arcsec) but seeing and sky brightness a bit variable. EL~67 at start.

02:35 Check open-loop tracking performance
Continue on Garnavich field, at least the guided data should be usable. Offset between pairs for dithering. Take 90s (30s x 3) with LUCIFER with guiding on and another 90s with guiding off at the same position.
  • +20" Y Offset 0162 guide-on 0163 guide-off
  • +20" X offset 0164 guide-on 0165 guide-off
  • -40" X offset 0166 guide-on 0167 guide-off
  • +20" Y offset 0168 guide-on 0169 guide-off
  • +20" X offset 0170 guide-on 0171 guide-off

03:10 Make new collimation model, implement, test
PSF arbitration set to "primary" for collimation model and M2 cleared to stay at its nominal global offset position. Luci files 0174-0191 minus some glitches. New collimation models made with and without "optimization" on. Currently running with the optimized one, but we suspect it will not be good with a large temperature change.

05:35 Take Field aberration measurements Running /home/LBTO/iraf/field_BS9107.cl in /OldRepository/AGW_Data/100103_BS9107. Doug watching the data.

*05:45 John's field_BS9107.cl script failed. Dave says IRAF is now in a funky state (must type all letter twice to see them...). Restarts IRAF. Doug check John's python scripts. Dave changes focal = "LDG" to "LFBG" in agw_collimate.py and agw_wfsc_multiple.py.

*05:59 Doug takes a wfsc image with IDL. Works fine wfscimage00001.fits

*06:02 Dave starts the field_BS9107.cl script. Doug sees data and all looks good.

*06:25 The telescope is not collimated and forces on the primary are getting high enough that new active optics corrections are not applied. Doug finds that his IDL routines are rotating the wrong direction to get into the primary coordinated system. He modified agw_collimage.pro agw_wfsc_multiple.pro and agw_post_collimate.pro. He collimated with IDL for three iteration (wfcs 25-27) and all is well.

*06:59 Dave restarts the IRAF script field_BS9107.cl. Doug monitors the collimate for a few on-axis sets and all looks good.

Guide Star wfscimage000*
0 28-31
82 32-34
0 35-36
35 37-39
0 40-41
19 42-44
0 45-46
64 47-49
0 50-51
96 52-54
0 55-56
40 57-59
0 60-61
35 62-64
0 65-66
48 67-69
0 70-71
8 72-74
0 75-76
68 77-79
0 80-81
96 82

Seeing got bad around files 75-76 and remained above 1.0 after that. Aborted after frame 82.

* 08:11 Restest updatePointingReference functionality.* Test needs to be done with IRTC commands as LUCIFER apparently cannot send the updatePointingReference command.
pointto BS9122 gs=0 and allow to collimate. EL~72.
pointto BS9122 gs=29 and test...
Set PSF arbitration back to "split" mode

LUCI# Position
0193 Preset position
0194 DETXY/REL +20 X
0195 updatePointingReference DETXY (did not move)
0196 DETXY/ABS 0,0 (did not move)
0197 DETXY/REL +20 X
0198 DETXY/ABS 0,0 (moves back to the updated PO

0199 DETXY/REL +20 X
0200 RADEC/ABS 0,0 (does not move - no RADEC offsets to absorb)

0--- RADEC/REL +20 N - EL axis went to HOLD mode, offset did not complete correctly.

08:37 preset again to BS9122 gs=29

0201 Preset position
0202 DETXY/REL -20 Y
0203 updatePointingReference DETXY (did not move)

0204 DETXY/REL -20 Y
0205 RADEC/REL +20 N
0206 RADEC/ABS 0,0 (goes to position -20 Y of current PO)

0207 updatePointingReference DETXY (did not move)
0208 updatePointingReference RADEC (did not move)

0209 DETXY/REL -20 Y
0210 RADEC/REL +20 N
0211 DETXY/ABS 0,0 (goes to position +20 N of current PO)
0212 RADEC/ABS 0,0 (goes to PO defined in 207/208)

0--- RADEC/REL +20 N - No Guide star found in guide box

09:10 preset again to BS9122 gs=29

0213 at preset position
0214 RADEC/REL +20 N
0215 updatePointingReference RADEC (did not move)
0216 RADEC/ABS 0,0 (did not move)
0217 RADEC/REL +20 N - No Guide star found in guide box, took image anyway because telescope appears to be tracking OK.

I need to think about this repeated loss of guide star, so going on to pointing.

09:24 Make new pointing model
pointing log started: /home/telescope/20100103.ptlog

UT PtStar Luci#   UT PtStar Luci#   UT PtStar Luci#   UT PtStar Luci#
09:53 WT10_403 0221 10:24 WT10_311 0234 10:55 WT10_0409 0247 11:20 WT10_163 0260
09:56 WT10_350 0222 10:26 WT10_243 0235 10:57 ACT0350 0248 11:23 WT10_165 0261
09:59 ACT0402 0223 10:28 ACT0307 0236 10:59 ACT0376 0249 11:25 ACT0184 0262
10:01 WT10_359 0224 10:30 ACT0186 0237 11:01 WT10_378 0250 11:26 ACT0239 0263
10:03 ACT0282 0225 10:33 ACT0086 0238 11:03 ACT0401 0251 11:28 WT10_186 0264
10:05 WT10_180 0226 10:35 WT10_131 0239 11:05 WT10_407 0252 11:32 ACT0152 0265
10:10 ACT0254 0227 10:37 WT10_273 0240 11:06 ACT0361 0253 11:33 WT10_120 0266
10:12 WT10_069 0228 10:40 WT10_252 0241 11:09 WT10_377 0254 11:35 WT10_190 0267
10:15 WT10_169 0229 10:42 ACT0214 0242 11:11 WT10_357 0255 11:37 WT10_243 0268
10:16 WT10_275 0230 10:45 ACT0301 0243 11:13 ACT0359 0256 11:39 WT10_309 0269
10:18 WT10_226 0231 10:47 ACT0335 0244 11:15 WT10_320 0257 11:41 WT10_311 0270
10:20 ACT0326 0232 10:50 ACT0404 0245 11:17 WT10_288 0258 11:43 WT10_298 0271
10:22 WT10_339 0233 10:52 WT10_400 0246 11:19 WT10_216 0259 11:44 WT10_367 0272

11:58 Seeing is OK (0.8 arcsec on guider). Executing martini SDT imaging:
martini_ksimg.txt ~5 pixels (0.6 arcsec) FWHM
martini_himg.txt ~4 pixels (0.5 arcsec) FWHM
martini_jimg.txt ~ pixels (0.X arcsec) FWHM
no standards?

Antennae Ks image 0303-0319 10sec x 6 integrated (two manual plus jitter 15 90.0)
Antenna sky 0320-0324 10s x 6

Arc measurement of rotator center on LUCIFER. All are 15x2s integrated
luci# PA
325 141 326 111 327 081 328 141 329 171 330 201

Darks in the morning...

Observing Plan

* X Correct initial collimation/pointing with IE/CA, test ptcorr script

* X Run short dither dice5 pattern(s) to check that things still work as in May/June

* Restest updatePointingReference functionality, notify LUCIFER Team of results.

* X Check open-loop tracking performance (rms on off-axis GS vs. rms on LUCIFER) as reference data for comparison to post-Nov-HBS work and this has not been checked in a while.

* Collect astrometric field data to check telescope scale at colder temps (use science data for this)

* X Check guide star FWHM/rms vs LUCIFER image quality (closed loop DIQ)

* X Make new collimation model, implement (optional if it looks good)
  • Derive and implement before pointing model obs

* Make new pointing model, implement (if coll model good)
  • Reset mount encoders (now takes only 2 minutes) @AZ=120, EL=90
  • Have TO start the pointing log!
  • Check first position in AZ < 0 degrees wrap
  • Use ARC method to measure AGw/Rotator center
  • 28 points are sufficient, more is always better
  • Can use NB filter in LUCIFER, the pointing stars are ~7.5 mag
  • needs stable temps!
  • May need to do several short PMs as temps are changing rapidly

* X Take Field aberration measurements if seeing is good (time-series, stability measurements)

* Demonstrate guiding to 18th mag (Tel=GUIDE)
* WFS on 18th mag. Will likely need to set the wfs max exposure time to 120s
* Quantify the guide/wfs limits with mag and seeing vs. LUCIFER DIQ

-- DavidThompson - 02 Jan 2010
Topic attachments
I Attachment Action Size Date Who Comment
daq.20100103.jpgjpg daq.20100103.jpg manage 677 K 07 Jan 2010 - 00:50 JohnHill Plot of Telescope and Mirror Temperatures
pmcc.sx.thermal.20100103.jpgjpg pmcc.sx.thermal.20100103.jpg manage 477 K 07 Jan 2010 - 00:51 JohnHill Plot of SX Primary Mirror Temperatures
Topic revision: r10 - 07 Jan 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback